OpenStudy (anonymous):

You have 7 old books and need to choose 4 to give away. How many different groups of 4 books can you make?

OpenStudy (anonymous):

I get 7*6*5=210, am I close?

OpenStudy (anonymous):

7c4 =7*6*4*3 = 504

OpenStudy (anonymous):

is a combinations problem, nCr in the calc if i remember right,

OpenStudy (anonymous):

I don't understand Raheen's answer - where is the 5?

OpenStudy (anonymous):

7c4 gives 35 in my calc...

OpenStudy (anonymous):

there is the combinations formula attached

OpenStudy (anonymous):

Thanks!

OpenStudy (luffingsails):

Hi Jenn, Raheen is not quite right. The key is that once you have chosen a book to give away... you can't choose it again! So, this is a pick without replacement problem. So, N-choose-R will work with some modification. The equation for N-choose-R without replacement is: N-Choose-R = N!/((R!)(N-R)! For this problem, that gives: 7!/((4!)(7-4)!) = 7!/(4!3!) = 35 Hope this helps.

OpenStudy (anonymous):

Yes! Thanks!